# Sequential estimation of disturbed aerodynamic flows from sparse measurements via a reduced latent space

### v1

@@ 3.2 Sequential filtering examples
Figure 14: The first, second, third, and eleventh observation modes at four distinct time instances during the gust interaction for the case ($\alpha=60^{\circ}$, $D_{y}=-0.71$, $\sigma=0.12$, $y_{o}=-0.19$, $t_{o}=3.3$). How the sensor numbers are ordered in the eigenvector plots is shown on the airfoil.
@@ 3.3 Sensor failure study
As shown in the preceding analysis, leading-edge sensors play a dominant role in informing the data assimilation process. A natural set of questions arises: What happens if a highly informative sensor fails or is removed? How significantly would this affect the accuracy of the state estimation? Would the estimator become unreliable simply due to the loss of a single, influential sensor? To address these questions, we investigate the effect of sensor dropout on flow estimation. A sensor can be effectively “dropped” by assigning it a much larger measurement noise variance, thereby diminishing its contribution in the assimilation update. Specifically, we simulate the failure of sensor 6, a leading-edge sensor identified as highly informative in figure 14, by increasing its variance by a factor of 100 relative to all other sensors. Data assimilation is then performed for the same flow condition presented in figure 14. For comparison, we also drop sensor 2, located near the trailing edge (TE) on the suction side—an area associated with minimal sensor response—in a separate setting and assess its impact relative to a baseline case where all sensors have equal uncertainty (see figures 10, 12, and 14).
@@
The results shown in figure 15 reveal that dropping either sensor leads to only a modest increase in the average lift prediction error, indicating that the estimator remains robust in capturing global aerodynamic loads. However, the impact on the reconstructed vorticity field is more pronounced, particularly when sensor 6—located near the leading edge (LE) on the pressure side—is removed. In this case, more noticeable degradation is observed in regions near the gust core and the downstream wake. The corresponding posterior uncertainty fields reveal a clear increase in the variance of the vorticity estimates in those regions, indicating that sensor 6 plays a critical role in constraining the latent state components associated with the downstream evolution of the gust-induced flow structures. In contrast, dropping sensor 2—located near the TE on the suction side—has a negligible effect on both the reconstructed flow field and the associated uncertainty bounds. This suggests that sensor 2 contributes minimally to the information content of the observation operator in this particular flow regime and sensor configuration. On the other hand, the removal of sensor 6, which had a strong contribution to the leading observation modes (as seen in figure 14), significantly alters the structure of the dominant observation eigenvectors, evident in figure 16. Specifically, the eigenmodes reconfigure by suppressing the weight of sensor 6 to nearly zero and compensating by reallocating higher weights to its neighboring sensors (e.g., sensors 5, 7, and 8). This redistribution allows the estimator to partially recover the lost observability; however, the reconfigured observation space may not fully span the same latent directions as before. As a result, the effective observation rank may decrease, or the information associated with critical latent directions may be weakened. This potential degradation in observability leads to increased posterior uncertainty along certain state directions, particularly those associated with the spatial regions that sensor 6 was originally sensitive to. This explains the elevated uncertainty and local deterioration in vorticity reconstruction quality observed in the gust-impacted regions when sensor 6 is dropped.
To further assess the estimator’s robustness, we examine the scenario in which multiple sensors, here sensors 2, 6, and 11, fail simultaneously. The corresponding results are shown in figures 15 and 16. To recall, sensor 6 originally provided the most informative measurements, so the overall reconstruction quality is comparable to the case where only sensor 6 is removed, evident in the dominant observation modes illustrated in figure 16. Interestingly, the vorticity reconstruction is slightly improved in the three-sensor failure case, which may be attributed to sensors 2 and 11 being highly correlated with other active sensors. Their removal can, at certain instants, reduce redundancy in the observation space and improve numerical conditioning, thereby marginally enhancing the filter’s performance in vorticity predictions. Interestingly, despite the simultaneous failure of 3 out of 11 sparse sensors (representing a $27\%$ outage), the estimator maintains robust performance.
Overall, by examining how dominant observation modes evolve under sensor dropout, one can identify the relative importance of each sensor, assess redundancy in the sensing architecture, and quantify the robustness of the estimation framework to partial sensor failure. The findings confirm that the current approach maintains resilience to non-critical sensor loss, with minimal degradation in flow estimation accuracy. Even in the case of dropping a highly informative sensor—such as sensor 6 near the LE—the estimator adapts by reallocating observational weight to neighboring sensors and continues to perform well in predicting global quantities like lift and capturing the dominant large-scale vortical structures in the flow. Although posterior uncertainty increases and local flow features near the gust core and wake become less precisely reconstructed, the overall aerodynamic response remains accurately captured. This highlights the robustness of the learned filtering framework and its potential for real-time deployment in scenarios with limited or partially compromised sensing.
@@
Figure 15: Effect of individual sensor dropout on the reconstructed lift (first panel) and vorticity fields (second to fourth panels) for the case ($\alpha=60^{\circ}$, $D_{y}=-0.71$, $\sigma=0.12$, $y_{o}=-0.19$, $t_{o}=3.3$).
@@
Figure 16: Effect of individual sensor dropout on the observation modes consistent with figure 15 at four representative time instances during the gust interaction. The condition is ($\alpha=60^{\circ}$, $D_{y}=-0.71$, $\sigma=0.12$, $y_{o}=-0.19$, $t_{o}=3.3$).
@@ 3.4 Sequential estimation for an extrapolated case
All training and test aerodynamic cases discussed thus far involved disturbances introduced during the first vortex shedding cycle. To evaluate the robustness and generalizability of the estimator, it is critical to test its performance under extrapolative conditions—specifically when a disturbance occurs at an unseen phase of the shedding cycle. For this purpose, we construct a test case where a dipole-like gust is introduced at $t_{o}=7$, well into the second shedding cycle, at the maximum AoA considered in this study, $\alpha=60^{\circ}$. We refer to this as the extrapolation case. The results are presented in figure 17, showing the predicted lift coefficient and vorticity fields at three representative time instances during the gust–airfoil interaction. Before the gust enters the domain, the predicted lift trajectory matches the undisturbed periodic behavior with excellent accuracy, reflecting the estimator’s ability to maintain nominal performance under regular flow conditions. Remarkably, once the gust is introduced at $t_{o}=7$, the pressure sensors immediately detect the upstream disturbance, resulting in a clear deviation from the undisturbed lift trajectory. This reflects the pressure field’s sensitivity to flow perturbations, even before their direct interaction with the airfoil around $t=7.4$. Importantly, since the forecast operator was trained only on post-disturbance dynamics, it cannot anticipate the effect of an incoming gust at $t_{o}=7$. As such, the forecasted latent states $\boldsymbol{\xi}(t_{o}+\Delta t|t_{o})$ remain confined to the undisturbed latent manifold. The deviation from undisturbed behavior is therefore initiated entirely by the correction introduced in the analysis step of filtering, wherein the discrepancy between observed and predicted pressures triggers an adjustment to the predicted states, resulting in the updated posterior $\boldsymbol{\xi}(t_{o}+\Delta t|t_{o}+\Delta t)$.
The reconstructed vorticity field reflects this behavior. The posterior mean closely captures the dominant shear layers and large-scale vortices in regions near the sensors. The temporal evolution of the flow is accurately tracked, including the unsteady shedding patterns and the influence of the gust. Notably, the estimated uncertainty fields highlight elevated variance in key regions: the gust core, the near-wake shear layers, and along the outer boundaries of the separated flow structures. These regions coincide with areas less directly impactful on pressure observations, indicating that the filter correctly expresses confidence where data is informative and uncertainty where it is not. This extrapolation test confirms that the filtering framework generalizes well beyond its training window, relying on pressure-based corrections to capture flow disturbances introduced at unseen phases. The pressure sensors effectively detect and drive state corrections in real time, while the uncertainty quantification provides meaningful confidence estimates. This underscores the estimator’s reliability and adaptability in operational settings where disturbances may arise unpredictably.
@@
One striking and consistent phenomenon observed across all estimated cases in Figures 9, 10, and 17 (as mentioned earlier) is the transient lift amplification caused by the interaction between the positive lobe of the dipole gust and the primary negative LE vortex. This interaction locally increases the effective angle of attack and momentarily elevates the lift coefficient. Shortly thereafter, the negative lobe of the dipole impinges on the airfoil, leading to a sharp decrease in lift. Notably, the largest estimation error and the highest uncertainty magnitude consistently occur during the time of maximum lift deviation from the undisturbed trajectory—precisely when the dipole is in closest interaction with the LE vortex. This phase corresponds to the most dynamically complex portion of the flow evolution, characterized by intense vortex interactions, rapid structural changes in the shear layer, and strong nonlinearity in the pressure response. The estimator reflects this complexity by exhibiting both increased prediction error and broader uncertainty bounds, indicating reduced observability and greater difficulty in capturing the rapidly changing flow state.
These demonstrations collectively highlight the capability of the proposed framework to transform sparse, noisy surface pressure measurements into reliable, uncertainty-aware reconstructions of high-dimensional flow fields in real time. By embedding neural-network surrogates within a low-rank ensemble filtering architecture, the method not only achieves computational efficiency but also adapts naturally to degraded sensing and extrapolative flow conditions. The result is an estimator that remains robust under severe disturbances, provides interpretable uncertainty quantification tied to observability, and scales efficiently for online deployment. Such capabilities establish a foundation for practical integration of flow state monitoring with feedback control and decision-making in unsteady aerodynamic environments.
@@ 4 Conclusion
This work proposed a data assimilation framework in a reduced space for real-time flow and load estimation in highly disturbed aerodynamic environments. By combining nonlinear autoencoders for compact state representation, neural-network-based surrogate models for dynamics and observations, and a low-rank Ensemble Kalman Filter (LREnKF), the proposed approach enables accurate, uncertainty-aware estimation of unsteady flow fields from sparse, noisy surface pressure measurements. The estimator was evaluated across a wide range of gust-disturbed aerodynamic scenarios, including both interpolative cases—where disturbances occurred during the first vortex shedding cycle, consistent with training—and extrapolative cases—where gusts were introduced at arbitrary instants across the shedding cycle. In all test scenarios, the predicted lift closely followed the true unsteady dynamics, and the reconstructed vorticity fields accurately captured the dominant shear layers and vortical structures, especially in regions adjacent to sensor locations. The estimator demonstrated strong generalization capabilities, reliably tracking complex flow behavior under conditions not encountered during training.
@@
To understand the limits of observability and estimation quality, the LREnKF was adopted to carry out state updates along the dominant directions of state and observation spaces. In all cases, the system was found to be effectively low-rank in the latent space, with only a few dominant latent directions being well-informed by the pressure observations. Similarly, a small number of observation directions—corresponding to linear combinations of sensor signals—were sufficient to capture most of the flow information. These findings highlight the underlying compression of the disturbed aerodynamic states and the selectivity of the sensors in constraining them. We further analyzed sensor informativeness over time by computing the dominant modes of the observation Gramian, which revealed the transient sensitivity of each pressure sensor to the evolving latent state. Leading-edge sensors were consistently the most informative, especially during gust–vortex interactions, while suction-side sensors near the trailing edge contributed less throughout the flow evolution. Sensor informativeness during gust encounters strongly depends on the gust structure. For example, in our previous work (Mousavi and Eldredge, 2025), a Taylor gust produced a more disruptive interaction, leading to different information content across sensors compared to the less disruptive and more compact Gaussian gust considered in this study. It is important to note that the LREnKF employed in this study assumes all sources of uncertainty to be Gaussian or well-approximated as such. Nonetheless, for strongly nonlinear and non-Gaussian systems, this framework can be extended using the gradient-based Gramian estimation approach proposed by Baptista et al., (2022).
In addition, we investigated the role of unobservable latent directions, i.e., directions in the latent space that lie in the nullspace of the gradient of the learned observation operator. Perturbations along these directions do not affect the predicted pressure observations but can still propagate through the decoder to influence the reconstructed vorticity field. This highlights a subtle but important limitation: although the filter cannot constrain these directions directly, their physical effect manifests in downstream flow quantities, potentially contributing to persistent uncertainty or bias in regions remote from or less informed by sensors. These unobservable latent directions also impact the estimator’s performance during the post-disturbance period, when the gust has convected downstream of the airfoil and the flow returns to its undisturbed state. In this interval, the estimator occasionally interprets residual uncertainties as spurious upstream gusts, resulting in artificially elevated uncertainty in regions where no actual disturbance exists.
@@
To probe the estimator’s robustness, we studied the impact of sensor dropout by artificially increasing the uncertainty of key sensors, effectively removing them from the assimilation process. Remarkably, even when the most informative leading-edge sensor was dropped, the estimator maintained high accuracy. This resilience stems from the low-rank structure of the system: when a dominant sensor is lost, the filter implicitly increases the Kalman gain weights on neighboring sensors, allowing them to compensate and reconstruct the dominant flow features. As a result, the degradation in lift and vorticity estimation was minimal, confirming that the estimator leverages spatial redundancy and shared observability among nearby sensors.
A key and consistent phenomenon observed across all test cases was the aerodynamic response to dipole gusts. When the positive lobe of the dipole interacted with the primary negative leading-edge vortex, it transiently increased the effective angle of attack, producing a sharp lift surge. This was followed by a rapid lift drop caused by the arrival of the negative lobe, which disrupted the vortex structure. These gust–vortex interactions corresponded to the most dynamically complex phases of the flow evolution and were consistently associated with the largest estimation errors and highest posterior uncertainty. The uncertainty fields correctly identified regions of limited observability—such as the gust core and wake shear layers—indicating that the estimator not only reconstructs the dominant flow features but also provides reliable confidence bounds.
@@
Overall, the results demonstrate that the proposed framework achieves accurate and efficient flow state estimation under strong disturbances, with robustness to sensor dropout and generalization to unseen gust realizations. By exploiting low-rank structure and integrating uncertainty quantification, this method provides a scalable and reliable foundation for flow monitoring, feedback control, and decision-making in real-time aerodynamic systems where traditional solvers are computationally prohibitive.

### v2

@@ 3.2 Sequential filtering examples
Figure 11: The first three observation modes at five distinct time instances during the gust-airfoil interaction for the case ($\alpha=60^{\circ}$, $D_{y}=-0.71$, $\sigma=0.12$, $y_{o}=-0.19$, $t_{o}=3.3$). How the sensor numbers are ordered in the eigenvector plots is shown on the airfoil.
The eigendeocmposition of the observation-space Gramian enables a systematic analysis of how and which measurement sensors contribute most significantly to the update step. It is thus insightful to examine the observation modes corresponding to the sensor configuration and their role in informing state estimation. Figure 11 presents the leading observation eigenmodes—specifically $u_{1}$, $u_{2}$, and $u_{3}$. The sensor indices follow the ordering shown in figure 2, where sensor 1 is located at the trailing edge on the upper surface, and the numbering proceeds counterclockwise around the airfoil, ending at the trailing edge on the lower surface. Each observation mode is a weighted combination of sensor outputs that, through the Kalman gain, most strongly influences the filter correction in a corresponding latent direction. At each time instant, the observation eigenmodes are ordered by descending eigenvalue magnitude, representing the most to least informative directions in the observation space. For instance, $u_{1}$ reflects the most informative observation direction. The weight assigned to each sensor in these modes reflects its relative contribution to that direction in the observation space.
From these plots, we observe that during the critical time interval $t\in[3.6,4.0]$, when the gust approaches and begins interacting with the airfoil on the pressure side, the leading-edge sensors 6 through 9 on the lower surface become the most responsive along $u_{1}$ direction. These sensors exhibit consistently high contributions across the dominant observation modes, indicating their central role in informing the estimator throughout the gust-airfoil-wake interaction ($t\in[3.6,5.0]$). While the relative contributions of sensors vary over time, the leading-edge and lower-surface sensors consistently remain the most informative along the dominant energetic mode. In contrast, the subsequent modes, $u_{2}$ and $u_{3}$, amplify the influence of the suction-side sensors, reflecting the deformation of the primary negative vortex induced by the gust. The temporal evolution of the modes may not appear entirely smooth at first glance. However, as the gust interacts with the airfoil, the eigenvalue spectrum of the observation Gramian evolves dynamically, occasionally leading to eigenvalue crossings and a subsequent reordering of the ranked modes. For instance, the mode labeled $u_{3}$ at $t=3.6$ may become mode $u_{2}$ at the next time $t=4.0$, while the previous $u_{2}$ shifts to a lower rank $u_{3}$. Or $u_{3}$ at $t=4.4$ becomes $u_{2}$ at $t=4.8$. This “mode reordering” has been consistently observed across all flow cases and reflects a physical reorganization of the sensor sensitivities—that is, a temporal shift in which sensor combinations dominate the information content.
It is worth emphasizing that this observation contrasts with findings from our earlier study [32], in which the upstream disturbance was modeled using a Taylor vortex. A Taylor vortex typically introduces a stronger and more disruptive disturbance compared to a vortex dipole, leading to intense interaction with the primary shear layer and edge vortices on the suction side. As a result, the suction-side sensors were found to be highly sensitive and informative in that setting. In contrast, the dipole-like gust generated by Gaussian forcing in the present work interacts more locally with the pressure side, producing a more focused and asymmetric sensor response. This distinction highlights the importance of gust structure in shaping the spatial distribution of sensor informativeness and ultimately determining the directions along which the latent states can be effectively constrained.
For the disturbed flow estimated in figure 9, the spectral decomposition of the state-space and observation-space Gramians at a representative time during the gust–airfoil interaction is shown in the left two panels of figure 12. The eigenvalues are arranged in descending order, corresponding to decreasing informativeness. The decay of the spectra indicates that the tail eigenmodes possess signal-to-noise ratios well below unity. These tail modes are therefore only weakly constrained by the pressure measurements, as the measurement-induced signals they generate lie within the uncertainty bounds of the sensors. Consequently, the state evolution along these weakly observed directions is governed primarily by the forecast propagation, with only negligible corrections introduced during the analysis step. Retaining $99\%$ of the cumulative spectral energy yields the effective state rank $r_{\boldsymbol{\xi}}$ and observation rank $r_{\boldsymbol{p}}$ shown in the right plot of figure 12. The state rank reflects the number of latent directions that are actively informed by the measurements and significantly corrected during assimilation, whereas the observation rank quantifies the number of independent pressure modes that effectively contribute to state updates. At early times, when the ensemble exhibits substantial deviation from the true state, corrections occur across all latent directions ($r_{\boldsymbol{\xi}}=7$). Once the state estimate converges toward the true trajectory, the number of dominant effective modes decreases and stabilizes, oscillating between four and five as the flow evolves.
Figure 12: The left two panels display the eigenvalue spectra of the state-space and observation-space Gramians at $t=2.4$. The right plot shows the corresponding effective state rank $r_{\boldsymbol{\xi}}$ and observation rank $r_{\boldsymbol{p}}$, defined as the minimum number of eigendirections required to retain $99\%$ of the cumulative spectral energy. The results are shown for the disturbed aerodynamic case with ($\alpha=20^{\circ}$, $D_{y}=-1.9$, $\sigma=0.1$, $y_{o}=0.22$, $t_{o}=1.5$).
Here, the terms “observed” and “weakly observed” are used to refer to latent directions that are strongly and weakly constrained by pressure measurements, respectively. The former correspond to modes with high observability, while the latter lie close to the approximate nullspace of the Jacobian of the observation operator, $\nabla\boldsymbol{h}_{\boldsymbol{W}}$, and therefore receive only minimal correction from the sensors. An important conclusion is that even the weakly observed tail modes are necessary to accurately reconstruct the vorticity field. Although corrections along weak modes are small at each individual assimilation step, their cumulative effect over time improves the latent state estimate. Since the decoder maps a low-dimensional latent vector to a high-dimensional vorticity field (on the order of $2.8\times 10^{4}$ degrees of freedom), even small latent corrections can yield noticeable improvements in the reconstructed flow.
For the case $\alpha=20^{\circ}$ under the same disturbed condition shown in figure 9, we further examine these weakly observed directions (corresponding to the smallest eigenvalues in figure 12) by explicitly constructing perturbations in the approximate nullspace of $\boldsymbol{C}_{\boldsymbol{p}}$ (which corresponds to the nullspace of $\nabla\boldsymbol{h}_{\boldsymbol{W}}$ in the deterministic limit). The resulting vorticity reconstructions, shown in figure 13, demonstrate that perturbations along these latent directions can produce predicted pressure measurements that remain within sensor uncertainty bounds, but still correspond to notable variation in the flow fields. This provides direct evidence of the many-to-one mapping between latent representations and pressure signatures. The induced vorticity differences are concentrated primarily in regions of limited sensor sensitivity, including the gust core, the leading-edge vicinity, the suction-side shear layer, and the wake. It confirms that weakly observed latent modes can contribute significantly to overall vorticity error. These results underscore a fundamental limitation of pressure-only sensing: in the absence of additional measurement modalities or prior regularization, multiple latent states may remain indistinguishable from the perspective of the sensors. This ambiguity becomes more pronounced after the gust convects downstream, when pressure sensitivity to residual flow variations decreases. Consequently, multiple plausible latent configurations may be indistinguishable from the available measurements during such regimes, and it increasingly falls on the forecast operator to propagate the latent state.
To mitigate this ambiguity, we explored incorporating a maximum-observability constraint during autoencoder training (section 2.2) to enforce a more uniquely pressure-consistent latent representation. While this approach improved latent state identifiability, it rendered the vorticity decoder highly sensitive to small latent perturbations, such that minor estimation errors produced non-physical artifacts and degraded reconstruction robustness. This tradeoff highlights the inherent tension between enhancing measurement-space observability and maintaining stability and robustness in high-dimensional flow reconstruction within compressed latent-variable frameworks.
Figure 13: The weakly-observed vorticity plots for disturbed flow at condition ($\alpha=20^{\circ}$, $D_{y}=-1.9$, $\sigma=0.1$, $y_{o}=0.22$, $t_{o}=1.5$).
@@ 3.3 Sensor failure study
As shown in the preceding analysis, leading-edge sensors (sensors 6 and 7) play a dominant role in informing the data assimilation process. A natural set of questions arises: What happens if a highly informative sensor fails or is removed? How significantly would this affect the accuracy of the state estimation? Would the estimator become unreliable simply due to the loss of a single, influential sensor? To address these questions, we investigate the effect of sensor dropout on flow estimation. A sensor can be effectively “dropped” by assigning it a much larger measurement noise variance, thereby diminishing its contribution in the assimilation update. Specifically, we simulate the failure of sensor 7, a leading-edge sensor identified as highly informative along $u_{1}$ in figure 11, by increasing its variance by a factor of 1000 relative to all other sensors. Data assimilation is then performed for the same flow condition presented in figure 11. For comparison, we also drop sensor 2, located near the trailing edge (TE) on the suction side—an area associated with minimal sensor response—in a separate setting and assess its impact relative to a baseline case where all sensors have equal uncertainty (see figures 10 and 11).
@@
The results shown in figure 14 reveal that dropping either sensor leads to only a modest increase in the average lift prediction error (this error is $\varepsilon=0.03$ for the case of no sensor being dropped), indicating that the estimator remains robust in capturing global aerodynamic loads. However, the impact on the reconstructed vorticity field is more pronounced, particularly when sensor 7—located near the leading edge (LE) on the pressure side—is removed. In this case, even with greater reconstruction error, the estimated vorticity field still captures the large-scale structures in the domain. In contrast, dropping sensor 2—located near the TE on the suction side—has a negligible effect on both the reconstructed flow field and the associated uncertainty bounds. This suggests that sensor 2 contributes minimally to the information content of the observation operator in this particular flow regime and sensor configuration. The removal of sensor 7, which had a strong contribution to the leading observation modes (as seen in figure 11), significantly alters the structure of the dominant eigenvectors of the observation Gramian $\boldsymbol{C}_{\boldsymbol{p}}$, evident in figure 15. Specifically, these observation eigenmodes reconfigure by suppressing the weight of sensor 7 to nearly zero and compensating by reallocating higher weights to its neighboring sensors. This redistribution allows the estimator to partially recover the lost observability; however, the reconfigured observation space may not fully span the same latent directions as before. It should be noted that although sensor 7 is highly informative in the leading observation eigenmode, other sensors carry comparable—sometimes even greater—information in the subsequent leading modes, as shown in figure 11, and thus can compensate the loss of sensor 7.
@@
Figure 14: Effect of individual sensor dropout on the reconstructed lift and vorticity fields for the case ($\alpha=60^{\circ}$, $D_{y}=-0.71$, $\sigma=0.12$, $y_{o}=-0.19$, $t_{o}=3.3$).
@@
Figure 15: Effect of individual sensor dropout on the observation modes consistent with figure 14 at four representative time instances during the gust-airfoil interaction. The condition is ($\alpha=60^{\circ}$, $D_{y}=-0.71$, $\sigma=0.12$, $y_{o}=-0.19$, $t_{o}=3.3$).
To further assess the estimator’s robustness, we examine the scenario in which multiple sensors, here sensors 2, 7, and 11, fail simultaneously. The corresponding results are shown in figures 14 and 15. To recall, sensor 7 originally provided the most informative measurements in the most energetic eigendirection $u_{1}$, so the overall reconstruction quality is comparable to the case where only sensor 7 is removed, evident in the dominant observation modes illustrated in figure 15. Interestingly, the vorticity reconstruction is slightly improved in the three-sensor failure case, which may be attributed to sensors 2 and 11 being highly correlated with other active sensors. Their removal can, at certain instants, reduce redundancy in the observation space and improve numerical conditioning, thereby marginally enhancing the filter’s performance in vorticity predictions. Interestingly, despite the simultaneous failure of 3 out of 11 sparse sensors (representing a $27\%$ outage), the estimator maintains robust performance.
Overall, by examining how dominant observation modes evolve under sensor dropout, one can identify the relative importance of each sensor, assess redundancy in the sensing architecture, and quantify the robustness of the estimation framework to partial sensor failure. The findings confirm that the current approach maintains resilience to sensor loss, with minimal degradation in flow estimation accuracy. Even in the case of dropping a highly informative sensor—such as sensor 7 near the LE—the estimator adapts by reallocating observational weight to neighboring sensors and continues to perform well in predicting global quantities like lift and capturing the dominant large-scale vortical structures in the flow. This highlights the robustness of the learned filtering framework and its potential for deployment in scenarios with limited or partially compromised sensing.
@@ 3.4 Sequential estimation for an extrapolated case
All training and test aerodynamic cases discussed thus far involved disturbances introduced during the first vortex shedding cycle. To evaluate the robustness and generalizability of the estimator, it is critical to test its performance under extrapolative conditions—specifically when a disturbance occurs at an unseen phase of the shedding cycle. For this purpose, we construct a test case where a dipole-like gust is introduced at $t_{o}=7$, well into the second shedding cycle, at the maximum AoA considered in this study, $\alpha=60^{\circ}$. We refer to this as the extrapolation case. The results are presented in figure 16, showing the predicted lift coefficient and vorticity fields at three representative time instances during the gust–airfoil interaction. Before the gust enters the domain, the predicted lift trajectory matches the undisturbed periodic behavior with excellent accuracy, reflecting the estimator’s ability to maintain nominal performance under regular flow conditions. When the gust is introduced at $t_{o}=7$, the pressure sensors immediately detect the upstream disturbance, resulting in a clear deviation from the undisturbed lift trajectory. This reflects the pressure field’s sensitivity to flow perturbations, even before their direct interaction with the airfoil around $t=7.4$. Importantly, the forecast operator is incapable of anticipating the effect of an incoming gust at $t_{o}=7$. As such, the forecasted latent states $\boldsymbol{\xi}(t_{o}+\Delta t|t_{o})$ remain confined to the undisturbed latent manifold. The deviation from undisturbed behavior is therefore initiated entirely by the correction introduced in the analysis step of filtering, wherein the discrepancy between observed and predicted pressures triggers an adjustment to the predicted states, resulting in the updated posterior $\boldsymbol{\xi}(t_{o}+\Delta t|t_{o}+\Delta t)$.
@@
The reconstructed vorticity field reflects this behavior. The posterior mean closely captures the dominant shear layers and large-scale vortices. The temporal evolution of the flow is accurately tracked, including the unsteady shedding patterns and the influence of the gust. Notably, the estimated uncertainty fields are so small that they are visually not detectable. This extrapolation test confirms that the filtering framework generalizes well beyond its training window, relying on pressure-based corrections to capture flow disturbances introduced at unseen phases. The pressure sensors effectively detect and drive state corrections quickly, while the uncertainty quantification provides meaningful confidence estimates. This underscores the estimator’s reliability and adaptability in operational settings where disturbances may arise unpredictably.
One striking and consistent phenomenon observed across all estimated cases in Figures 9, 10, and 16 (as mentioned earlier) is the transient lift spike caused by the interaction between the dipole gust and the boundary layers. Notably, the largest estimation error and the highest uncertainty magnitude consistently occur during the time of maximum lift deviation from the undisturbed trajectory—precisely when the dipole is in closest interaction with the primary edge vortex. This phase corresponds to the most dynamically complex portion of the flow evolution, characterized by intense vortex interactions, rapid structural changes in the shear layer, and strong nonlinearity in the pressure response. The estimator reflects this complexity by exhibiting both increased prediction error and broader uncertainty bounds, indicating reduced observability and greater difficulty in capturing the rapidly changing flow state.
Across all test cases, the estimator maintains accurate lift tracking and resolves the dominant vortical structures from sparse pressure measurements. By embedding NN surrogates within an ensemble filtering architecture, the method not only achieves computational efficiency but also adapts naturally to degraded sensing and extrapolative flow conditions. The result is an estimator that remains robust under severe disturbances, provides interpretable uncertainty quantification, and scales efficiently for online deployment. Such capabilities establish a foundation for practical integration of flow state monitoring with feedback control and decision-making in unsteady aerodynamic environments.
@@ 4 Conclusion
This work proposed a data assimilation framework in a reduced space for fast flow and load estimation in randomly disturbed aerodynamic environments. By combining nonlinear convolutional autoencoders for compact state representation, neural-network-based surrogate models for dynamics and observations, and ensemble Kalman filtering (EnKF), the proposed approach enables accurate, uncertainty-aware estimation of unsteady flow fields from sparse, noisy surface pressure measurements. It was discussed that the learned forecast operator does not capture the transition from undisturbed to disturbed trajectories without measurement correction. In the absence of assimilation, the forecast predicts that the latent state remains on the nominal baseline trajectory. The measurement update step introduces the deviation from this trajectory once the pressure signature of the disturbance appears, enabling the estimator to track the disturbed evolution. The estimator was evaluated across a wide range of gust-disturbed aerodynamic scenarios, including both interpolative cases—where disturbances occurred during the first vortex shedding cycle, consistent with training—and extrapolative cases—where gusts were introduced at arbitrary instants across the shedding cycle. In all test scenarios, the predicted lift closely followed the true unsteady dynamics, and the reconstructed vorticity fields accurately captured the dominant vortical structures, especially in regions adjacent to sensor locations. The estimator demonstrated strong generalization capabilities, reliably tracking complex flow behavior under conditions not encountered during training.
@@
To quantify the limits of observability, we performed eigenvalue decompositions of the state-space and observation-space Gramians. Across all cases, the system exhibits intrinsically weakly observed directions under surface pressure sensing. In the original vorticity space, these directions correspond primarily to the gust core and portions of the downstream wake. These regions were shown to lie approximately in the nullspace of the Jacobian of the observation operator, indicating that perturbations along these directions produce negligible pressure signatures. We further analyzed sensor informativeness over time by computing the dominant modes of the observation Gramian, which revealed the transient sensitivity of each pressure sensor to the evolving latent state. Leading-edge sensors were consistently the most informative in the first few energetic modes, especially during gust–vortex interactions, while suction-side sensors near the trailing edge contributed less throughout the flow evolution. Sensor informativeness during gust encounters strongly depends on the gust structure. For example, in our previous work [32], a Taylor gust produced a more disruptive interaction, leading to different information content across sensors compared to the less disruptive and more compact Gaussian gust considered in this study. It is important to note that the EnKF employed in this study assumes all sources of uncertainty to be Gaussian or well-approximated as such. Nonetheless, for strongly nonlinear and non-Gaussian systems, this framework can be extended using the gradient-based Gramian estimation approach proposed by Baptista et al. [3].
@@
To probe the estimator’s robustness, we studied the impact of sensor dropout by artificially increasing the uncertainty of key sensors, effectively removing them from the assimilation process. Remarkably, even when the most informative leading-edge sensor was dropped, the estimator maintained high accuracy. This resilience stems from the mode redistribution of the system: when a dominant sensor is lost, the filter implicitly increases the Kalman gain weights on neighboring sensors, allowing them to compensate and reconstruct the dominant flow features. As a result, the degradation in lift and vorticity estimation was minimal, confirming that the estimator leverages spatial redundancy and shared observability among nearby sensors.
@@
The gust–vortex interactions corresponded to the most dynamically complex phases of the flow evolution and were consistently associated with the largest estimation errors and highest posterior uncertainty. The uncertainty fields correctly identified regions of physically-limited observability—such as the gust core and wake shear layers—indicating that the estimator not only reconstructs the dominant flow features but also provides reliable confidence bounds. Overall, the results demonstrate that the proposed framework achieves accurate and efficient flow state estimation under strong arbitrary disturbances, with robustness to sensor dropout and generalization to unseen gust realizations. By exploiting sequential filtering, this method provides a scalable and reliable foundation for flow monitoring, feedback control, and decision-making in aerodynamic systems where traditional solvers are computationally prohibitive.